OpenStudy (anonymous):

Find the equation of line(s) tangent to y=(x-4)/(x-1) and parallel to y=3x-7. I know m=3 and f'(x)=-3/(x-1)^2 then it's set equal to 3 and after that I'm stumped. My book is not helping f'(x)=(1)(x-4)-(x-1)(1)/(x-1)^2 f'(x)=x-4-x+1/(x-1)^2 f'(x)=-3/(x-1)^2 so -3/(x-1)^2=3 then???? Stumped

OpenStudy (anonymous):

solve for x.

OpenStudy (jhannybean):

Set the derivative equal to the slope of the parallel line

OpenStudy (jhannybean):

First lets heck t see if your derivative is correct.

OpenStudy (anonymous):

f'(x)=3/(x-1)^2

OpenStudy (jhannybean):

alrighty..

OpenStudy (anonymous):

now proceed..equate the derivative to 3 and solve for x.

OpenStudy (anonymous):

I'm not clear on how to solve for x - my algebra is rusty

OpenStudy (jhannybean):

\[\large \frac{3}{(x-1)^2}=3\]\[\large 3=3(x-1)^2\]divide by 3 on both sides \[\large \frac{3}{3}=\frac{3(x-1)^2}{3}\]\[\large 1=(x-1)^2\]

OpenStudy (anonymous):